# PHENIX Collaboration

### v1

@@ Abstract
The PHENIX experiment has performed a systematic study of identified charged-hadron ($\pi^{\pm}$, $K^{\pm}$, $p$, $\bar{p}$) production at midrapidity in $p$$+$Al, ${}^{3}$e$+$Au, Cu$+$Au collisions at $\sqrt{s_{{}_{NN}}}=200$ GeV and U$+$U collisions at $\sqrt{s_{{}_{NN}}}=193$ GeV. Identified charged-hadron invariant transverse-momentum ($p_{T}$) and transverse-mass ($m_{T}$) spectra are presented and interpreted in terms of radially expanding thermalized systems. The particle ratios of $K/\pi$ and $p/\pi$ have been measured in different centrality ranges of large (Cu$+$Au, U$+$U) and small ($p$$+$Al, ${}^{3}$e$+$Au) collision systems. The values of $K/\pi$ ratios measured in all considered collision systems were found to be consistent with those measured in $p$$+$$p$ collisions. However the values of $p/\pi$ ratios measured in large collision systems reach the values of $\approx 0.6$, which is $\approx 2$ times larger than in $p$$+$$p$ collisions. These results can be qualitatively understood in terms of the baryon enhancement expected from hadronization by recombination. Identified charged-hadron nuclear-modification factors ($R_{AB}$) are also presented. Enhancement of proton $R_{AB}$ values over meson $R_{AB}$ values was observed in central ${}^{3}$e$+$Au, Cu$+$Au, and U$+$U collisions. The proton $R_{AB}$ values measured in $p$$+$Al collision system were found to be consistent with $R_{AB}$ values of $\phi$, $\pi^{\pm}$, $K^{\pm}$, and $\pi^{0}$ mesons, which may indicate that the size of the system produced in $p$$+$Al collisions is too small for recombination to cause a noticeable increase in proton production.
@@ I INTRODUCTION
Small collision systems (such as $p$$+$Al, ${}^{3}$e$+$Au) were previously studied primarily to investigate cold-nuclear-matter effects [25, 26] (Cronin enhancement [27, 28], multiple-parton scattering [29], modifications of the initial nuclear-parton distribution functions [30, 31]). It was thought [32, 25] that the size and lifetime of the system produced in small collision systems are not sufficient for QGP formation. However, in 2019 the PHENIX experiment reported on the observation of elliptic and triangular flow patterns of charged particles produced in $p$/$d$/${}^{3}$e$+$Au collisions [32]. The results of these measurements have been described in the frame of hydrodynamical models, which include the formation of short-lived QGP droplets. Comparison of identified charged-hadron production in small and large collision systems may enable systematic studies to determine the minimal conditions of QGP formation and to investigate influence of collision geometry and system size on hadron production.
This paper presents new data on identified charged-hadron production in $p$$+$Al, ${}^{3}$e$+$Au, Cu$+$Au collisions at $\sqrt{s_{{}_{NN}}}$ = 200 GeV and U$+$U collisions at $\sqrt{s_{{}_{NN}}}$ = 193 GeV. Discussed are the influence of collision centrality and geometry on identified charged-hadron invariant $p_{T}$ and $m_{T}$ spectra, nuclear-modification factors, and ratios.
@@ II DATA ANALYSIS
The data sets used in the present analysis were collected by the PHENIX experiment during the operational periods of the Relativistic Heavy Ion Collider in calendar years 2012 (Cu$+$Au and U$+$U), 2014 (${}^{3}$e$+$Au), and 2015 ($p$$+$Al). The PHENIX experiment [22] comprised several subdetectors that are grouped into four main blocks called arms. The central arms—east and west—cover midrapidity ($|\eta|<0.35$) and are intended for measurement of electrons, photons, and charged hadrons. In the present analysis, the drift chambers (DC) [33] are used for momentum determination, the time-of-flight (TOF) wall [34] is used for charged hadron identification, and the beam-beam counters (BBC) [35] are used for centrality determination. Detailed information about the PHENIX experiment can be found in Refs. [22, 36, 35].
@@ II.2 Particle Identification
where $p_{0}$, $p_{1}$, $p_{2}$, $p_{3}$, and $p_{4}$ are fit parameters, and the set of fit parameters is different for $m_{\rm TOF}^{2}(p_{T})$ and $\sigma_{\rm TOF}(p_{T})$.
@@ II.3.1 Reconstruction Efficiency
Figure 2: $\delta_{\rm FD}$ values as a function of $p_{T}$ for protons and antiprotons, measured in $p$$+$Al, ${}^{3}$e$+$Au, Cu$+$Au, and U$+$U collisions. The shaded bands indicate the systematic uncertainty.
The geometric acceptance of the detectors, detector efficiency, and analysis cuts have been taken into account by the reconstruction efficiency factor ($\epsilon_{\rm rec}$), estimated using a single-particle Monte-Carlo (MC) simulation. The MC simulation was carried out using the PHENIX integrated simulation application (PISA) package [38], which is based on GEANT3 [39] simulation software. In the PISA project the geometry, material, spatial, momentum and energy resolutions of the PHENIX detector subsystems as well as the configuration of the magnetic field are simulated in full accordance with the structure of the real facility. Simulated events have been processed using the same analysis criteria as the data. In that way, the reconstruction efficiency $\epsilon_{\rm rec}$($p_{T}$) can be estimated by the number of reconstructed hadrons ($N_{\rm rec}^{\rm MC}$) divided by the total number of hadrons generated in the MC model ($N_{\rm tot}^{\rm MC}$):
@@
II.3.2 Weak Decay Correction
@@
MC simulations of $\Lambda$ and $\bar{\Lambda}$ are used to calculate the fraction of $p$ and $\bar{p}$ that come from $\Lambda$ and $\bar{\Lambda}$ decays. Figure 2 shows the values of $\delta_{\rm FD}$ as a function of $p_{T}$, obtained in $p$$+$Al, ${}^{3}$e$+$Au, Cu$+$Au, and U$+$U collisions.
@@ II.5 Invariant Spectra
The $\pi^{\pm}$, $K^{\pm}$, $p$, and $\bar{p}$ invariant transverse momentum spectra were measured in $p$$+$Al, ${}^{3}$e$+$Au, Cu$+$Au collisions at $\sqrt{s_{{}_{NN}}}$ = 200 GeV and in U$+$U collisions at $\sqrt{s_{{}_{NN}}}$ = 193 GeV as follows:
@@
where $T$ is the inverse-slope parameter and $A$ is a normalization factor. Examples of resulting $\pi^{+}$, $K^{+}$ and $p$ invariant-$m_{T}$ spectra measured in central Cu$+$Au collisions are presented in Fig. 5. Approximations by Eq. 5 are shown with [red] solid lines. Values of inverse-slope parameters $T$ calculated in $p$$+$Al, ${}^{3}$e$+$Au, Cu$+$Au, and U$+$U collisions are summarized in Table 6.
@@ III RESULTS
Figure 3: The $\pi^{+}$, $\pi^{-}$, $K^{+}$,$K^{-}$, $p$ and $\bar{p}$ invariant $p_{T}$ spectra measured in different centralities of $p$$+$Al, ${}^{3}$e$+$Au, Cu$+$Au collisions at $\sqrt{s_{{}_{NN}}}$ = 200 GeV and U$+$U collisions at $\sqrt{s_{{}_{NN}}}$ = 193 GeV. Invariant $p_{T}$ spectra are multiplied by powers of ten for clarity of presentation.
@@
Figure 6: Freeze-out temperature ($T_{0}$) as a function of $\langle N_{\rm part}\rangle$ obtained for positively and negatively charged hadrons at different centralities of $p$$+$Al, ${}^{3}$e$+$Au, Cu$+$Au, and U$+$U collisions. The $T_{0}$ values measured in $p$$+$$p$ collisions are shown for comparison.
Figure 6: Freeze-out temperature ($T_{0}$) as a function of $\langle N_{\rm part}\rangle$ obtained for positively and negatively charged hadrons at different centralities of $p$$+$Al, ${}^{3}$e$+$Au, Cu$+$Au, and U$+$U collisions. The $T_{0}$ values measured in $p$$+$$p$ collisions are shown for comparison.
Figure 7: Collective velocities ($u_{t}\right>$) as a function of $\langle N_{\rm part}\rangle$ obtained for positively and negatively charged hadrons in different centralities of $p$$+$Al, ${}^{3}$e$+$Au, Cu$+$Au, and U$+$U collisions. The $u_{t}\right>$ values measured in $p$$+$$p$ collisions are shown for comparison.
Table 5: Identified charged hadron inverse slope parameters $T$ (MeV/$c^{2}$) calculated in $p$$+$Al, ${}^{3}$e$+$Au, Cu$+$Au, U$+$U collision systems.
Table 6: Freeze-out temperatures and averaged collective velocities calculated for positively charged ($T_{0}^{+}$, $u_{t}\right>^{+}$) and negatively charged ($T_{0}^{-}$, $u_{t}\right>^{-}$) hadrons at different centralities of $p$$+$Al, ${}^{3}$e$+$Au, Cu$+$Au, and U$+$U collisions
@@
. System $\langle N_{\rm part}\rangle$ $T_{0}^{+}$ $T_{0}^{-}$ $u_{t}\right>^{+}$ $u_{t}\right>^{-}$ $p$$+$Al 3.1$\pm$0.1 167.99 $\pm$ 2.17 166.49 $\pm$ 15.29 0.28 $\pm$ 0.01 0.32 $\pm$ 0.04 4.4$\pm$0.3 171.84 $\pm$ 0.92 171.33 $\pm$ 15.96 0.29 $\pm$ 0.01 0.32 $\pm$ 0.04 3.3$\pm$0.1 168.01 $\pm$ 2.15 167.88 $\pm$ 14.30 0.28 $\pm$ 0.01 0.31 $\pm$ 0.04 1.6$\pm$0.2 164.10 $\pm$ 4.39 163.80 $\pm$ 11.20 0.27 $\pm$ 0.01 0.30 $\pm$ 0.03 ${}^{3}$e$+$Au 11.3$\pm$0.5 189.16 $\pm$ 10.22 166.27 $\pm$ 1.70 0.33 $\pm$ 0.03 0.38 $\pm$ 0.01 21.1$\pm$1.3 193.84 $\pm$ 12.42 166.44 $\pm$ 3.23 0.34 $\pm$ 0.03 0.39 $\pm$ 0.01 15.4$\pm$0.9 188.99 $\pm$ 9.82 167.59 $\pm$ 1.24 0.33 $\pm$ 0.02 0.38 $\pm$ 0.01 9.5$\pm$0.6 185.05 $\pm$ 8.57 164.60 $\pm$ 0.27 0.31 $\pm$ 0.02 0.36 $\pm$ 0.01 4.8$\pm$0.3 177.02 $\pm$ 4.99 158.33 $\pm$ 3.45 0.28 $\pm$ 0.01 0.33 $\pm$ 0.01 Cu$+$Au 70.4$\pm$3.0 154.06 $\pm$ 16.63 176.03 $\pm$ 11.95 0.46 $\pm$ 0.03 0.41 $\pm$ 0.02 154.8$\pm$4.1 149.99 $\pm$ 23.98 172.82 $\pm$ 19.75 0.48 $\pm$ 0.05 0.44 $\pm$ 0.04 80.4$\pm$3.3 157.02 $\pm$ 14.55 178.11 $\pm$ 10.20 0.45 $\pm$ 0.03 0.40 $\pm$ 0.02 34.9$\pm$2.9 161.02 $\pm$ 3.86 177.79 $\pm$ 2.35 0.40 $\pm$ 0.01 0.36 $\pm$ 0.01 11.5$\pm$1.8 160.04 $\pm$ 4.50 175.86 $\pm$ 7.56 0.34 $\pm$ 0.01 0.30 $\pm$ 0.02 U$+$U 330$\pm$6 160.03 $\pm$ 12.02 170.80 $\pm$ 13.24 0.48 $\pm$ 0.02 0.46 $\pm$ 0.03 259$\pm$7 169.04 $\pm$ 0.60 174.76 $\pm$ 2.93 0.44 $\pm$ 0.01 0.43 $\pm$ 0.01 65$\pm$6 176.03 $\pm$ 11.45 182.61 $\pm$ 9.64 0.39 $\pm$ 0.02 0.37 $\pm$ 0.02
Table 6: Freeze-out temperatures and averaged collective velocities calculated for positively charged ($T_{0}^{+}$, $u_{t}\right>^{+}$) and negatively charged ($T_{0}^{-}$, $u_{t}\right>^{-}$) hadrons at different centralities of $p$$+$Al, ${}^{3}$e$+$Au, Cu$+$Au, and U$+$U collisions
The dotted lines on Fig. 5 represent linear fits of the $T(m_{0})$ values from each centrality bin by Eq. 6. The fit parameters for positively charged ($T_{0}^{+}$, $u_{t}\right>^{+}$) and negatively charged ($T_{0}^{-}$, $u_{t}\right>^{-}$) hadrons calculated in $p$$+$Al, ${}^{3}$e$+$Au, Cu$+$Au, and U$+$U collision systems are shown in the Table 6 and presented in Figs. 7 and 7 as a function of $\langle N_{\rm part}\rangle$ values. Presented error values for $T_{0}^{\pm}$ and $u_{t}\right>^{\pm}$ parameters are due to the fit uncertainties. The $T_{0}$ values calculated in collisions with different geometries and centralities were found to be coincident within uncertainties, indicating that the freeze-out temperature is approximately independent of $\langle N_{\rm part}\rangle$ values. In contrast, the strength of the average transverse flow ($u_{t}\right>$ values) increases from small to large $\langle N_{\rm part}\rangle$ values, supporting the hydrodynamic picture.
Figure 8: The ratios of $p$/$\pi^{+}$ as a function of $p_{T}$ measured in different centralities of ${}^{3}$e$+$Au collisions and $\bar{p}$/$\pi^{-}$ ratios measured in different centralities of $p$$+$Al and ${}^{3}$e$+$Au collisions. Data points, measured in $p$$+$$p$ collisions [43], are shown as a comparison.
Figure 8: The ratios of $p$/$\pi^{+}$ as a function of $p_{T}$ measured in different centralities of ${}^{3}$e$+$Au collisions and $\bar{p}$/$\pi^{-}$ ratios measured in different centralities of $p$$+$Al and ${}^{3}$e$+$Au collisions. Data points, measured in $p$$+$$p$ collisions [43], are shown as a comparison.
@@
Figure 10: The ratios of $K^{+}$/$\pi^{+}$ and $K^{-}$/$\pi^{-}$ as a function of $p_{T}$ measured in different centralities of $p$$+$Al and ${}^{3}$e$+$Au collisions. Data points measured in $p$$+$$p$ collisions [43] are shown for comparison.
Figure 10: The ratios of $K^{+}$/$\pi^{+}$ and $K^{-}$/$\pi^{-}$ as a function of $p_{T}$ measured in different centralities of $p$$+$Al and ${}^{3}$e$+$Au collisions. Data points measured in $p$$+$$p$ collisions [43] are shown for comparison.
@@
Figure 12: Identified charged hadron nuclear-modification factors as a function of $p_{T}$ measured in central and peripheral $p$$+$Al, $d$$+$Au and ${}^{3}$e$+$Au collisions. The dashed lines are drawn as a visual aid at the value of $R_{AB}$ = 1 indicating absence of nuclear modification.
Figure 12: Identified charged hadron nuclear-modification factors as a function of $p_{T}$ measured in central and peripheral $p$$+$Al, $d$$+$Au and ${}^{3}$e$+$Au collisions. The dashed lines are drawn as a visual aid at the value of $R_{AB}$ = 1 indicating absence of nuclear modification.
@@
Figure 15: Comparison of $\phi$ [46, 26], $\pi^{\pm}$, $K^{\pm}$, $(p+\bar{p})/2$ and $\pi^{0}$ [47] $R_{AB}$ values vs. $p_{T}$ measured in central and peripheral $p$$+A$l and ${}^{3}$e$+$Au collisions. The dashed lines are drawn as a visual aid at the value of $R_{AB}$ = 1 indicating absence of nuclear modification.

### v2

@@ Abstract
The PHENIX experiment has performed a systematic study of identified charged-hadron ($\pi^{\pm}$, $K^{\pm}$, $p$, $\bar{p}$) production at midrapidity in $p$$+$Al, 3He$+$Au, Cu$+$Au collisions at $\sqrt{s_{{}_{NN}}}=200$ GeV and U$+$U collisions at $\sqrt{s_{{}_{NN}}}=193$ GeV. Identified charged-hadron invariant transverse-momentum ($p_{T}$) and transverse-mass ($m_{T}$) spectra are presented and interpreted in terms of radially expanding thermalized systems. The particle ratios of $K/\pi$ and $p/\pi$ have been measured in different centrality ranges of large (Cu$+$Au, U$+$U) and small ($p$$+$Al, 3He$+$Au) collision systems. The values of $K/\pi$ ratios measured in all considered collision systems were found to be consistent with those measured in $p$$+$$p$ collisions. However the values of $p/\pi$ ratios measured in large collision systems reach the values of  $\approx\,0.6$, which is  $\approx\,2$ times larger than in $p$$+$$p$ collisions. These results can be qualitatively understood in terms of the baryon enhancement expected from hadronization by recombination. Identified charged-hadron nuclear-modification factors ($R_{AB}$) are also presented. Enhancement of proton $R_{AB}$ values over meson $R_{AB}$ values was observed in central 3He$+$Au, Cu$+$Au, and U$+$U collisions. The proton $R_{AB}$ values measured in $p$$+$Al collision system were found to be consistent with $R_{AB}$ values of $\phi$, $\pi^{\pm}$, $K^{\pm}$, and $\pi^{0}$ mesons, which may indicate that the size of the system produced in $p$$+$Al collisions is too small for recombination to cause a noticeable increase in proton production.
@@ I INTRODUCTION
Small collision systems (such as $p$$+$Al, 3He$+$Au) were previously studied primarily to investigate cold-nuclear-matter effects [25, 26] (Cronin enhancement [27, 28], multiple-parton scattering [29], modifications of the initial nuclear-parton distribution functions [30, 31]). It was thought [32, 25] that the size and lifetime of the system produced in small collision systems are not sufficient for QGP formation. However, in 2019 the PHENIX experiment reported on the observation of elliptic and triangular flow patterns of charged particles produced in $p$/$d$/3He$+$Au collisions [32]. The results of these measurements have been described in the frame of hydrodynamical models, which include the formation of short-lived QGP droplets. Comparison of identified charged-hadron production in small and large collision systems may enable systematic studies to determine the minimal conditions of QGP formation and to investigate influence of collision geometry and system size on hadron production.
This paper presents new data on identified charged-hadron production in $p$$+$Al, 3He$+$Au, Cu$+$Au collisions at $\sqrt{s_{{}_{NN}}}$ = 200 GeV and U$+$U collisions at $\sqrt{s_{{}_{NN}}}$ = 193 GeV. Discussed are the influence of collision centrality and geometry on identified charged-hadron invariant $p_{T}$ and $m_{T}$ spectra, nuclear-modification factors, and ratios.
@@ II DATA ANALYSIS
The data sets used in the present analysis were collected by the PHENIX experiment during the operational periods of the Relativistic Heavy Ion Collider in calendar years 2012 (Cu$+$Au and U$+$U), 2014 (3He$+$Au), and 2015 ($p$$+$Al). The PHENIX experiment [22] comprised several subdetectors that are grouped into four main blocks called arms. The central arms—east and west—cover midrapidity ($|\eta|<0.35$) and are intended for measurement of electrons, photons, and charged hadrons. In the present analysis, the drift chambers (DC) [33] are used for momentum determination, the time-of-flight (TOF) wall [34] is used for charged hadron identification, and the beam-beam counters (BBC) [35] are used for centrality determination. Detailed information about the PHENIX experiment can be found in Refs. [22, 36, 35].
@@ II.2 Particle identification
where $p_{0}$, $p_{1}$, $p_{2}$, $p_{3}$, and $p_{4}$ are fit parameters, and the set of fit parameters is different for $m_{\rm TOF}^{2}(p_{T})$ and $\sigma_{\rm TOF}(p_{T})$. In the high-$p_{T}$ region, hadron signals start to overlap; therefore, the veto cut was introduced for better separation of $\pi$, $K$ and $p$. The veto cut requires that the hadron mass does not satisfy the $1.5\,\sigma$ condition for neighboring hadrons. The PID and veto cuts are standard for the PHENIX detector [14, 17].
@@ II.3.1 Reconstruction efficiency
Figure 2: $\delta_{\rm FD}$ values as a function of $p_{T}$ for protons and antiprotons, measured in $p$$+$Al, 3He$+$Au, Cu$+$Au, and U$+$U collisions. The shaded bands indicate the systematic uncertainty.
The geometric acceptance of the detectors, detector efficiency, and analysis cuts have been taken into account by the reconstruction efficiency factor ($\epsilon_{\rm rec}$), estimated using a single-particle Monte-Carlo (MC) simulation. The MC simulation was carried out using the PHENIX integrated simulation application (PISA) package [38], which is based on geant[39] simulation software. In the PISA project the geometry, material, spatial, momentum and energy resolutions of the PHENIX detector subsystems as well as the configuration of the magnetic field are simulated in full accordance with the structure of the real facility. Simulated events have been processed using the same analysis criteria as the data. In that way, the reconstruction efficiency $\epsilon_{\rm rec}$($p_{T}$) can be estimated by the number of reconstructed hadrons ($N_{\rm rec}^{\rm MC}$) divided by the total number of hadrons generated in the MC model ($N_{\rm tot}^{\rm MC}$):
@@
II.3.2 Weak-decay correction
@@
MC simulations of $\Lambda$ and $\bar{\Lambda}$ are used to calculate the fraction of $p$ and $\bar{p}$ that come from $\Lambda$ and $\bar{\Lambda}$ decays. Figure 2 shows the values of $\delta_{\rm FD}$ as a function of $p_{T}$, obtained in $p$$+$Al, 3He$+$Au, Cu$+$Au, and U$+$U collisions.
@@ II.5 Invariant spectra
The $\pi^{\pm}$, $K^{\pm}$, $p$, and $\bar{p}$ invariant transverse momentum spectra were measured in $p$$+$Al, 3He$+$Au, Cu$+$Au collisions at $\sqrt{s_{{}_{NN}}}$ = 200 GeV and in U$+$U collisions at $\sqrt{s_{{}_{NN}}}$ = 193 GeV as follows:
@@
where $T$ is the inverse-slope parameter and $A$ is a normalization factor. Examples of resulting $\pi^{+}$, $K^{+}$ and $p$ invariant-$m_{T}$ spectra measured in central Cu$+$Au collisions are presented in Fig. 5. Approximations by Eq. 5 are shown with [red] solid lines. Values of inverse-slope parameters $T$ calculated in $p$$+$Al, 3He$+$Au, Cu$+$Au, and U$+$U collisions are summarized in Table 6.
@@ III RESULTS
Figure 3: The $\pi^{+}$, $\pi^{-}$, $K^{+}$,$K^{-}$, $p$ and $\bar{p}$ invariant $p_{T}$ spectra measured in different centralities of $p$$+$Al, 3He$+$Au, Cu$+$Au collisions at $\sqrt{s_{{}_{NN}}}$ = 200 GeV and U$+$U collisions at $\sqrt{s_{{}_{NN}}}$ = 193 GeV. Invariant $p_{T}$ spectra are multiplied by powers of ten for clarity of presentation.
@@
Figure 6: Freeze-out temperature ($T_{0}$) as a function of $\langle N_{\rm part}\rangle$ obtained for positively and negatively charged hadrons at different centralities of $p$$+$Al, 3He$+$Au, Cu$+$Au, and U$+$U collisions. The $T_{0}$ values measured in $p$$+$$p$ collisions are shown for comparison.
Figure 7: Collective velocities ($u_{t}\right>$) as a function of $\langle N_{\rm part}\rangle$ obtained for positively and negatively charged hadrons in different centralities of $p$$+$Al, 3He$+$Au, Cu$+$Au, and U$+$U collisions. The $u_{t}\right>$ values measured in $p$$+$$p$ collisions are shown for comparison.
Table 5: Identified charged hadron inverse slope parameters $T$ (MeV/$c^{2}$) calculated in $p$$+$Al, 3He$+$Au, Cu$+$Au, U$+$U collision systems.
@@
Table 6: Freeze-out temperatures and averaged collective velocities calculated for positively charged ($T_{0}^{+}$, $u_{t}\right>^{+}$) and negatively charged ($T_{0}^{-}$, $u_{t}\right>^{-}$) hadrons at different centralities of $p$$+$Al, 3He$+$Au, Cu$+$Au, and U$+$U collisions
. System $\langle N_{\rm part}\rangle$ $T_{0}^{+}$ $T_{0}^{-}$ $u_{t}\right>^{+}$ $u_{t}\right>^{-}$ $p$$+$Al 3.1$\pm$0.1 167.99 $\pm$ 2.17 166.49 $\pm$ 15.29 0.28 $\pm$ 0.01 0.32 $\pm$ 0.04 4.4$\pm$0.3 171.84 $\pm$ 0.92 171.33 $\pm$ 15.96 0.29 $\pm$ 0.01 0.32 $\pm$ 0.04 3.3$\pm$0.1 168.01 $\pm$ 2.15 167.88 $\pm$ 14.30 0.28 $\pm$ 0.01 0.31 $\pm$ 0.04 1.6$\pm$0.2 164.10 $\pm$ 4.39 163.80 $\pm$ 11.20 0.27 $\pm$ 0.01 0.30 $\pm$ 0.03 3He$+$Au 11.3$\pm$0.5 189.16 $\pm$ 10.22 166.27 $\pm$ 1.70 0.33 $\pm$ 0.03 0.38 $\pm$ 0.01 21.1$\pm$1.3 193.84 $\pm$ 12.42 166.44 $\pm$ 3.23 0.34 $\pm$ 0.03 0.39 $\pm$ 0.01 15.4$\pm$0.9 188.99 $\pm$ 9.82 167.59 $\pm$ 1.24 0.33 $\pm$ 0.02 0.38 $\pm$ 0.01 9.5$\pm$0.6 185.05 $\pm$ 8.57 164.60 $\pm$ 0.27 0.31 $\pm$ 0.02 0.36 $\pm$ 0.01 4.8$\pm$0.3 177.02 $\pm$ 4.99 158.33 $\pm$ 3.45 0.28 $\pm$ 0.01 0.33 $\pm$ 0.01 Cu$+$Au 70.4$\pm$3.0 154.06 $\pm$ 16.63 176.03 $\pm$ 11.95 0.46 $\pm$ 0.03 0.41 $\pm$ 0.02 154.8$\pm$4.1 149.99 $\pm$ 23.98 172.82 $\pm$ 19.75 0.48 $\pm$ 0.05 0.44 $\pm$ 0.04 80.4$\pm$3.3 157.02 $\pm$ 14.55 178.11 $\pm$ 10.20 0.45 $\pm$ 0.03 0.40 $\pm$ 0.02 34.9$\pm$2.9 161.02 $\pm$ 3.86 177.79 $\pm$ 2.35 0.40 $\pm$ 0.01 0.36 $\pm$ 0.01 11.5$\pm$1.8 160.04 $\pm$ 4.50 175.86 $\pm$ 7.56 0.34 $\pm$ 0.01 0.30 $\pm$ 0.02 U$+$U 330$\pm$6 160.03 $\pm$ 12.02 170.80 $\pm$ 13.24 0.48 $\pm$ 0.02 0.46 $\pm$ 0.03 259$\pm$7 169.04 $\pm$ 0.60 174.76 $\pm$ 2.93 0.44 $\pm$ 0.01 0.43 $\pm$ 0.01 65$\pm$6 176.03 $\pm$ 11.45 182.61 $\pm$ 9.64 0.39 $\pm$ 0.02 0.37 $\pm$ 0.02
[…]
Figure 8: The ratios of $p$/$\pi^{+}$ as a function of $p_{T}$ measured in different centralities of 3He$+$Au collisions and $\bar{p}$/$\pi^{-}$ ratios measured in different centralities of $p$$+$Al and 3He$+$Au collisions. Data points, measured in $p$$+$$p$ collisions [43], are shown as a comparison.
@@
Figure 10: The ratios of $K^{+}$/$\pi^{+}$ and $K^{-}$/$\pi^{-}$ as a function of $p_{T}$ measured in different centralities of $p$$+$Al and 3He$+$Au collisions. Data points measured in $p$$+$$p$ collisions [43] are shown for comparison.
@@
Figure 12: Identified charged hadron nuclear-modification factors as a function of $p_{T}$ measured in central and peripheral $p$$+$Al, $d$$+$Au and 3He$+$Au collisions. The dashed lines are drawn as a visual aid at the value of $R_{AB}$ = 1 indicating absence of nuclear modification.
@@
Figure 15: Comparison of $\phi$ [46, 26], $\pi^{\pm}$, $K^{\pm}$, $(p+\bar{p})/2$ and $\pi^{0}$ [47] $R_{AB}$ values vs. $p_{T}$ measured in central and peripheral $p$$+A$l and 3He$+$Au collisions. The dashed lines are drawn as a visual aid at the value of $R_{AB}$ = 1 indicating absence of nuclear modification.